# Partonic distribution functions and amplitudes using tensor network methods

### v1

@@ Abstract
Calculations of the parton distribution function (PDF) and distribution amplitude (DA) are highly relevant to core experimental programs as they provide non-perturbative inputs to inclusive and exclusive processes, respectively. Direct computation of the PDFs and DAs remains challenging because they are non-perturbative quantities defined as light-cone correlators of quark and gluon fields, and are therefore inherently time-dependent. In this work, we propose a uniform quantum strategy based on tensor network simulation techniques to directly extract these hadronic quantities from first principles using the matrix product state of the hadrons in the same setup. We present exemplary numerical calculations with the Nambu-Jona-Lasinio model in 1+1 dimensions and compare with available exact diagonalization and quantum circuit simulation results. Using tensor networks, we evaluate the PDF and DA at various strong couplings in the large-qubit limit, which is consistent with expectations at perturbative and non-relativistic limits.
@@ 1 Introduction
Quantum computing technology has undergone rapid development and may provide an alternative path toward calculating the structure of hadrons from first principles Bauer:2022hpo ; DiMeglio:2023nsa . The exponential Hilbert space and quantum simulation algorithms provide a natural and efficient theoretical framework to simulate quantum field theory Jordan:2011ci ; Jordan:2014tma ; Jordan:2017lea . Studying real-time evolution. such as the case for the PDF and DA, is particularly suited to quantum simulation Lamm:2019uyc . The idea is to encode the fermionic and gluonic degrees of freedom into spin Hamiltonians and simulate the real-time matrix elements for evaluating these distribution functions and amplitudes directly on the quantum circuit. Recently, exploratory studies using quantum computing for parton structures are booming: from low-dimensional (1+1)-dimensional field theories Lamm:2019uyc ; Perez-Salinas:2020nem ; LiTianyin:2021kcs ; Li:2022lyt ; LiTianyin:2024nod ; Banuls:2024oxa ; Barata:2024bzk ; Grieninger:2024cdl ; Grieninger:2024axp , worldline formulation Mueller:2019qqj , spacetime Wilson loop Echevarria:2020wct , to effective light-front Hamiltonian models Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp , these studies already exhibit interesting physics and potential. However, most quantum simulation for parton structure is still restricted in performing systematic calculation and limited in extending to the continuum, largely due to today’s noisy intermeidate-scale quantum infrastructure Preskill:2018jim .
@@
In this work, we aim to leverage tensor network (TN) methods in computing hadron structures. Traditionally used in condensed matter physics, TN have become an increasingly promising tool of its own to study hadronic physics and high-energy physics Silvi:2017srb ; Banuls:2018jag . Owing to its similarity with quantum computing, TN, despite being a classical computation approach, is used as a natural testbed for quantum simulation, especially for low dimensional and low entanglement problems. In particular, TN is an efficient simulation strategy to simulate quantum-many body systems by representing the quantum states locally in the Hilbert space and their dynamics can often be bounded by an area-law of entanglement entropy Orus:2013kga ; mps-reps-gs . Importantly, TN is free from sign problems and reliable in accessing regimes inaccessible to Monte Carlo methods Banuls:2018jag . Here, we aim to compute both the PDF and DA in the same setting of the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions using efficient TN simulation algorithms. Having direct access to hundreds of qubits with TN, we present numerical simulation at large-qubit limit that is representative of the continuum limit. Specifically, we compare the behavior of the PDF and DA at various mass values and strong coupling strengths to make qualitative comparison with available results in perturbative QCD and the non-relativistic limits.
@@ 2.2 Nambu-Jona-Lasinio model in 1+1 dimensions
Now we compute the PDF in eq. (1) and DA in eq. (3) using the obtained (1+1)-dimensional NJL Hamiltonian. Specifically, we use $\gamma^{+}=\frac{1}{\sqrt{2}}(\gamma^{0}+\gamma^{1})=\frac{1}{\sqrt{2}}(\sigma^{z}-i\sigma^{y})$ and $\vec{n}=(1,-1)$ for the light-cone coordinate. It follows that $\vec{n}\cdot{\vec{\gamma}}=\sqrt{2}\gamma^{+}$ and $\vec{n}t=(1,-1)z$ with a reparameterized variable $z$ for space and time. Since the gauge fields are absent in the NJL model, the explicit Wilson lines in the PDF and DA vanish and we are left with time evolution operators of the Hamiltonians. While this feature of the NJL model dramatically simplifies the calculation, we are well aware that one should include dynamical gluon contributions in a more realistic scenario with available methods for simulating the Wilson line Zohar:2012ts ; Brennen:2015pgn ; Lamm:2018siq ; Zohar:2019tsa ; Echevarria:2020wct , which we leave for a subsequent investigation. Applying the same JW encoding to field operators $\psi^{\dagger}$ and $\psi$, the final expression of the PDF and DA are related to the sum of four matrix element expectations,
@@ 3 Results
In this work, we prepare the hadron state and simulate the real-time correlators of the quark and gluon fields using tensor network framework ITensor ITensor:2022 . The numerical results are compared with available exact diagonalization and quantum simulation results. Throughout this paper, we always use the lattice spacing $a=1$ in our simulations.
@@ 3.3 Parton distribution function and amplitudes
To demonstrate numerical results in the continuum limit, we use sufficiently high number of qubits in our simulations to compute the PDF and DA using tensor networks. Such calculation is beyond the current limits of exact simulation and quantum simulation with classical hardware. With tensor networks, we are able to study physical results at continuum limits utilizing an efficient representation of the many-body system with the MPS. In Fig. 2, we compute both the PDF and DA in the same setting with increasing system sizes (i.e., qubits) using selected mass $m=0.7a^{-1}$ and strong coupling $g=0.25$. For the PDFs, as expected in a one-flavor hadron (such as $\pi^{0}$, composed of many pairs of $q$ and $\bar{q}$), we see an asymmetric plot with a peak at $x=0.5$ and $x=-0.5$. Similarly for the DAs, we see a typical distribution centered in the $x\in[0,1]$ region. As system size $N$ increases, both the PDF and DA results gradually converge to the large-qubit curve at $N=102$, suggesting the calculations approach the continuum of $N\rightarrow\infty$ limit [^4] In principle, one would take both $a\rightarrow 0$ and $N\rightarrow\infty$. For simplicity, we fix $a=1$.. Therefore, for all the results below, we always use $N=102$ qubits for the simulation and we find it sufficient to represent the continuum behavior of the PDF and DA at this large qubit limit for our NJL Hamiltonian. Overall, the shape and tendency of our results are qualitatively consistent with experimental observations such as E791 data E791:2000xcx for the PDF and Fermilab E615 data Conway:1989fs for the DA of the pion.
@@ 4 Conclusion and outlook
In this work, we proposed a quantum strategy based on tensor network simulation to extract the PDF and DA for hadrons in the Nambu-Jona-Lasinio model in 1+1 dimensions. Our approach utilizes the DMRG algorithm to accurately construct the hadron and vacuum states on a charge-preserving lattice and the TDVP algorithm to efficiently time evolve the two-point correlator needed for these non-perturbative quantities. With our approach, we extract the PDF and DA in the same setting with converging results up to 102 qubits. We are able to study the behavior of the PDF and DA under the same setting for various masses and strong couplings, and observe consistent agreement with the expectations from the perturbative QCD as well as non-relativistic limit. Our exploratory study shows the advantage of tensor network in the current era for computing various physically important quantities in the continuum limit.
It would be very important to expand non-perturbative calculations of quantities such as the PDF, DA, and hadronic tensors using tensor network to U(1) and SU(2) gauge theories with consideration of multiple spins and flavors. Inclusion of flavor degrees of freedom in the line of study would provide significant understanding to realistic mesons and baryons from first principles. While tensor networks have primarily been applied to 1+1 dimensions, it would be interesting to extend to higher lattice dimensions. The natural extension of MPS to higher dimensions would be projected entanglement pair states Verstraete:2004cf ; verstraete2004valence , tree tensor network Shi:2005ctp , and so forth, which have been used to study 3+1 dimension Lattice Gauge Theory Magnifico:2020bqt ; Banuls:2018jag . We also plan to incorporate more advanced simulation techniques, such as tensor renormalization group LevinNave:2007 and its natural extension, tensor network renormalization tnr , to scale to higher qubit simulations with reduced error and study entanglement structure. These methods can be used to calculate correlation functions and generalize to systems beyond one spatial dimension.

### v2

@@ Abstract
Calculations of the parton distribution function (PDF) and distribution amplitude (DA) are highly relevant to core experimental programs as they provide non-perturbative inputs to inclusive and exclusive processes, respectively. Direct computation of the PDFs and DAs remains challenging because they are non-perturbative quantities defined as light-cone correlators of quark and gluon fields, and are therefore inherently time-dependent. In this work, we use a uniform quantum strategy based on tensor network simulation techniques to directly extract these hadronic quantities from first principles using the matrix product state of the hadrons in the same setup. We present exemplary numerical calculations with the Nambu-Jona-Lasinio model in 1+1 dimensions and compare with available exact diagonalization and quantum circuit simulation results. Using tensor networks, we evaluate the PDF and DA at various strong couplings in the large-qubit limit, which is consistent with expectations at perturbative and non-relativistic limits.
@@ 1 Introduction
Quantum computing technology has undergone rapid development and may provide an alternative path toward calculating the structure of hadrons from first principles Bauer:2022hpo ; DiMeglio:2023nsa . The exponential Hilbert space and quantum simulation algorithms provide a natural and efficient theoretical framework to simulate quantum field theory Jordan:2011ci ; Jordan:2014tma ; Jordan:2017lea . Studying real-time evolution. such as the case for the PDF and DA, is particularly suited to quantum simulation Lamm:2019uyc . The idea is to encode the fermionic and gluonic degrees of freedom into spin Hamiltonians and simulate the real-time matrix elements for evaluating these distribution functions and amplitudes directly on the quantum circuit. Recently, exploratory studies using quantum computing for parton structures are booming: from low-dimensional (1+1)-dimensional field theories Lamm:2019uyc ; Perez-Salinas:2020nem ; LiTianyin:2021kcs ; Li:2022lyt ; LiTianyin:2024nod ; Grieninger:2024cdl ; Grieninger:2024axp , worldline formulation Mueller:2019qqj , spacetime Wilson loop Echevarria:2020wct , to effective light-front Hamiltonian models Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp , these studies already exhibit interesting physics and potential. However, most quantum simulation for parton structure is still restricted in performing systematic calculation and limited in extending to the continuum, largely due to today’s noisy intermeidate-scale quantum infrastructure Preskill:2018jim .
@@
In this work, we aim to leverage tensor network (TN) methods in computing hadron structures. Traditionally used in condensed matter physics, TN have become an increasingly promising tool of its own to study hadronic physics and high-energy physics Silvi:2017srb ; Banuls:2018jag . Owing to its similarity with quantum computing, TN, despite being a classical computation approach, is used as a natural testbed for quantum simulation, especially for low dimensional and low entanglement problems. In particular, TN is an efficient simulation strategy to simulate quantum-many body systems by representing the quantum states locally in the Hilbert space and their dynamics can often be bounded by an area-law of entanglement entropy Orus:2013kga ; mps-reps-gs . Importantly, TN is free from sign problems and reliable in accessing regimes inaccessible to Monte Carlo methods Banuls:2018jag . Recently, TN methods are successfully demonstrated in evaluations of hadronic vacuum polarization Barata:2024bzk and the PDF Banuls:2024oxa in the 1+1 QED for the first time. Here, we aim to compute both the PDF and DA in the same setting of the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions using efficient TN simulation algorithms. Having direct access to hundreds of qubits with TN, we present numerical simulation at large-qubit limit that is representative of the continuum limit. Specifically, we compare the behavior of the PDF and DA at various mass values and strong coupling strengths to make qualitative comparison with available results in perturbative QCD and the non-relativistic limits.
@@ 2.2 Nambu-Jona-Lasinio model in 1+1 dimensions
Now we compute the PDF in eq. (1) and DA in eq. (3) using the obtained (1+1)-dimensional NJL Hamiltonian. Specifically, we use $\gamma^{+}=\frac{1}{\sqrt{2}}(\gamma^{0}+\gamma^{1})=\frac{1}{\sqrt{2}}(\sigma^{z}-i\sigma^{y})$ and $\vec{n}=(1,-1)$ for the light-cone coordinate. It follows that $\vec{n}\cdot{\vec{\gamma}}=\sqrt{2}\gamma^{+}$ and $\vec{n}t=(1,-1)z$ with a reparameterized variable $z$ for space and time. Since the gauge fields are absent in the NJL model, the explicit Wilson lines in the PDF and DA vanish and we are left with time evolution operators of the Hamiltonians. While this feature of the NJL model dramatically simplifies the calculation, we are well aware that one should include dynamical gluon contributions in a more realistic scenario with available methods for simulating the Wilson line Zohar:2012ts ; Brennen:2015pgn ; Lamm:2018siq ; Zohar:2019tsa ; Echevarria:2020wct , which we leave for a subsequent investigation. Recently, treatment of evolution of the Wilson lines was proposed and detailed for the Schwinger model Banuls:2024oxa . Applying the same JW encoding to field operators $\psi^{\dagger}$ and $\psi$, the final expression of the PDF and DA are related to the sum of four matrix element expectations,
@@ 3 Results
In this work, we prepare the hadron state and simulate the real-time correlators of the quark and gluon fields using tensor network framework ITensor ITensor:2022 . The numerical results are compared with available exact diagonalization and quantum simulation results. Throughout this paper, we always use the lattice spacing $a=1$ in our simulations. Importantly, our simulation follows the implementation using DMRG for state preparation and the TDVP for time evolution; whereas the recent work on Schwinger model uses DMRG and TEBD respectively Banuls:2024oxa .
@@ 3.3 Parton distribution function and amplitudes
To demonstrate numerical results in the continuum limit, we use sufficiently high number of qubits in our simulations to compute the PDF and DA using tensor networks. Such calculation is beyond the current limits of exact simulation and quantum simulation with classical hardware. With tensor networks, we are able to study physical results at continuum limits utilizing an efficient representation of the many-body system with the MPS. In Fig. 2, we compute both the PDF and DA in the same setting with increasing system sizes (i.e., qubits) using selected mass $m=0.7a^{-1}$ and strong coupling $g=0.25$. For the PDFs, as expected in a one-flavor hadron (such as $\pi^{0}$, composed of many pairs of $q$ and $\bar{q}$), we see an asymmetric plot with a peak at $x=0.5$ and $x=-0.5$. Similarly for the DAs, we see a typical distribution centered in the $x\in[0,1]$ region. As system size $N$ increases, both the PDF and DA results gradually converge to the large-qubit curve at $N=102$, suggesting the calculations approach the continuum of $N\rightarrow\infty$ limit [^4] In principle, one would take both $a\rightarrow 0$ and $N\rightarrow\infty$. For simplicity, we fix $a=1$.. Therefore, for all the results below, we always use $N=102$ qubits for the simulation and we find it sufficient to represent the continuum behavior of the PDF and DA at this large qubit limit for our NJL Hamiltonian. Overall, the shape and tendency of our results are qualitatively consistent with experimental observations such as E791 data E791:2000xcx for the PDF and Fermilab E615 data Conway:1989fs for the DA of the pion and they also resemble numerical calculations in 1+1 QCD Jia:2018qee .
@@ 4 Conclusion and outlook
In this work, we used a quantum strategy based on tensor network simulation to extract the PDF and DA for hadrons in the Nambu-Jona-Lasinio model in 1+1 dimensions in the same setting. Our approach utilizes the DMRG algorithm to accurately construct the hadron and vacuum states on a charge-preserving lattice and the TDVP algorithm to efficiently time evolve the two-point correlator needed for these non-perturbative quantities. With our approach, we extract the PDF and DA in the same setting with converging results up to 102 qubits. We are able to study the behavior of the PDF and DA under the same setting for various masses and strong couplings, and observe consistent agreement with the expectations from the perturbative QCD as well as non-relativistic limit. Our exploratory study shows the advantage of tensor network in the current era for computing various physically important quantities in the continuum limit.
It would be very important to expand non-perturbative calculations of quantities such as the PDF, DA, and hadronic tensors using tensor network to U(1) and SU(2) gauge theories. Notably, similar calculation using TN methods has been successfully demonstrated in Ref. Banuls:2024oxa for the U(1) case in the Schwinger model. Inclusion of multiple flavor and spin degrees of freedom in the line of study would provide significant understanding to realistic mesons and baryons from first principles. While tensor networks have primarily been applied to 1+1 dimensions, it would be interesting to extend to higher lattice dimensions. The natural extension of MPS to higher dimensions would be projected entanglement pair states Verstraete:2004cf ; verstraete2004valence , tree tensor network Shi:2005ctp , and so forth, which have been used to study 3+1 dimension Lattice Gauge Theory Magnifico:2020bqt ; Banuls:2018jag . We also plan to incorporate more advanced simulation techniques, such as tensor renormalization group LevinNave:2007 and its natural extension, tensor network renormalization tnr , to scale to higher qubit simulations with reduced error and study entanglement structure. These methods can be used to calculate correlation functions and generalize to systems beyond one spatial dimension.